OpenStudy (anonymous):

I don't seem to be able to do this problem without looking at the answer lol \[\int sin^{3}xcos^{2}x dx\] do I try to eliminate either sin or cos?

OpenStudy (australopithecus):

Integral of sin^(3)(x)(1-sin^(2))

OpenStudy (australopithecus):

Now expand

OpenStudy (australopithecus):

integral of sin^(3)(x)dx - integral of sin^(5)(x)dx

OpenStudy (anonymous):

\[\int sin^{3}(x)(1-sin^{2})\]

OpenStudy (australopithecus):

now you have two integrals

OpenStudy (anonymous):

\[\int sin^{3}(x)dx - \int sin^{5}(x)dx\]

OpenStudy (australopithecus):

use substitution

OpenStudy (fwizbang):

Its better to go the other way \[\int\limits_{?}^{?}\sin ^3 x \cos ^2 x = \int\limits_{?}^{?} sinx \cos^2 x ( 1 - \cos ^2x) dx\] and then switch variables to u =cos x.

OpenStudy (australopithecus):

\[\int\limits_{}^{}\sin^{2}(x)\sin(x)dx - \int\limits_{}^{}\sin(x)(\sin^{2})^{2}dx\]

OpenStudy (australopithecus):

\[\int\limits_{}^{}(1-\cos^{2}(x))\sin(x)dx - \int\limits_{}^{}\sin(x)(1-\cos^{2}(x))^{2}\]

OpenStudy (australopithecus):

you can just keep splitting them up like this till the integral is solved

OpenStudy (australopithecus):

\[\int\limits_{}^{}\sin(x)dx - \int\limits_{}^{} \cos^{2}(x)\sin(x)dx + \int\limits_{}^{}\sin(x) - \int\limits_{}^{}\sin(x)(1-\cos^{2}(x))^{2}dx\]

OpenStudy (australopithecus):

your prof must really hate you lol

OpenStudy (anonymous):

<falls off chair> there must be an easier way

OpenStudy (australopithecus):

soo tedious

OpenStudy (anonymous):

I think I found something here I just can't get it in order

OpenStudy (australopithecus):

here is the instructions for dealing with integrals such as these provided in my calc 1 class

OpenStudy (australopithecus):

you have to do it this way

OpenStudy (australopithecus):

or at least this is the only way I know how to do it

OpenStudy (australopithecus):

here are*

OpenStudy (anonymous):

u=cosx du=-sinxdx \[\int sin^{2}xcos^{2}xsinxdx \]

OpenStudy (australopithecus):

I would suggest doing the long simple method :) this isnt so bad eventually you will have the answer

OpenStudy (anonymous):

You shouldnt be able to solve it like that lol. sin3cos2 = sin 3 - sin 5 disregard the first term sin5 = sin2sin3 = sin3 - sin3cos2 youre back at the start ahaha.

OpenStudy (australopithecus):

just do it my way trust me

OpenStudy (anonymous):

\[\int (1-cos^{2})xcos^{2}xsinxdx\]

OpenStudy (anonymous):

\[\int (1-cos^{2}x)cos^{2}xsinxdx\]

OpenStudy (australopithecus):

integrals can be long and tedious but you have to deal with it, at least they dont require complex methods

OpenStudy (australopithecus):

well at least this one doesnt

OpenStudy (australopithecus):

some do

OpenStudy (anonymous):

u=cosx du=-sinxdx \[\int (1-cos^{2}x)cos^{2}xsinxdx\] \[-\int (1-u^{2})u^{2}du\]

OpenStudy (australopithecus):

the time you spent messing with other techniques you could have finished this integral lol

OpenStudy (fwizbang):

Once again Its better to go the other way \[ ∫sin^3xcos^2x=∫ sinxcos^2x(1−cos^2x)dx\] and then switch variables to u =cos x. du =-sin x dx so the integral becomes \[\int\limits_{?}^{?} (-du) u^2(1-u^2) = \int\limits_{?}^{?} du (u^4 -u^2)\]

OpenStudy (anonymous):

principle of nus tradamos

OpenStudy (anonymous):

sigh ok

OpenStudy (anonymous):

either way it's an integral, long and tidious

OpenStudy (anonymous):

thanks everyone!

OpenStudy (anonymous):

\[\frac{1}5 cos^{5}x -\frac 1 3 cos^{3} +C\]

OpenStudy (australopithecus):

its not so bad its very straight forward thus there isn't room for much mistake, also this document might help you through your adventures in the land of calc 1

OpenStudy (australopithecus):

it covers pretty much everything covered in calc 1 at my university very briefly

OpenStudy (australopithecus):

it is where that image came from

OpenStudy (anonymous):

i see, thanks @Australopithecus

OpenStudy (unklerhaukus):

Is this in your standrad table of integrals ? \[\int\sin^m(x)\cos^n(x)\text dx\]\[=\frac{\sin^{m+1}(x)\cos^{n-1}}{m+n}+\frac{n-1}{m+n}\int\sin(x)^m\cos^{n-2}(x)\text dx\]

OpenStudy (anonymous):

lets see \[\int\sin^3(x)\cos^2(x)\text dx\] \[=\frac{\sin^{4}(x)\cos(x)}{5}+\frac{1}{5}\int\sin(x)^3\text dx\]

OpenStudy (anonymous):

something seems off though

OpenStudy (anonymous):

the answer should be in cosines

OpenStudy (anonymous):

\[\frac 1 5 cos^ 5-\frac 1 3 cos^ 3 x + C\]

OpenStudy (unklerhaukus):

some trig substitution

OpenStudy (anonymous):

that would make the last term still \[\frac 1 {20} sin^4\]

OpenStudy (anonymous):

I would think?

OpenStudy (unklerhaukus):

hmm